OpenStudy (anonymous):

Write the equation y = 2 - 4x in standard form 4x - y = -2 4x + y = 2 4x + 2y = 0 y - 2 = 4(x - 0)

OpenStudy (anonymous):

what are the possibal answers?

OpenStudy (anonymous):

@erielle

OpenStudy (anonymous):

4x - y = -2 4x + y = 2 4x + 2y = 0 y - 2 = 4(x - 0

OpenStudy (anonymous):

4x+y=2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw